# Prethermalization, shadowing breakdown, and the absence of Trotterization transition in quantum circuits

### v1

@@ IV Conclusion
We have shown that the truncated propagator is a very versatile method to study Floquet systems. It primarily determines evolution of correlation functions, specifically, through the largest eigenvalue and the associated eigenvector it gives an explicit prediction for the prethermalization time and the effective (local) Floquet Hamiltonian governing prethermalization. This allowed us to resolve the question about reliability of approximating Hamiltonian evolution by a small-step Trotterization. Except for a single almost conserved observable all other orthogonal ones do not exhibit prethermalization, and even the almost conserved one decays in the thermodynamic limit. This means that due to the many-body dynamics the so-called shadowing time is always finite, and that there is no Trotterization transition in chaotic quantum systems. Previous observations were a consequence of finite-size effects as one has to ensure that the Heisenberg time is larger than the (long) prethermalization time.
@@
The truncated propagator can be also used to identify prethermal discrete time crystals. Just looking at the spectrum will already give an indication if there is a prethermal DTC (eigenvalues very close to roots of $1$, e.g., those around $-1$ in Fig. 2C at $\tau=3.0$). Studying chaotic kicked Ising model we find, perhaps surprisingly, that the prethermalized regime takes up the larger part of the gate duration interval $\tau\in[0,\pi]$. For instance, prethermalization is absent (say $\Delta>10^{-3}$, i.e. $T<10^{3}$) only for roughly $\tau\in[1.5,2.2]$. Glimpses of this rather slow thermalization in the kicked Ising model can be seen already in data presented in Ref. prosen07 . Prethermalization seems to be more common that previously thought and is not limited only to large frequencies (small $\tau$).
@@
Momentum dependence of the largest eigenvalue also contains information about transport. We have used it to calculate the energy diffusion constant during the prethermalization phase, as well as for the Hamiltonian model obtained in the limit $\tau\to 0$.
@@
While we always used numerics to obtain the spectrum of the truncated propagator, it would be interesting to see if exact solutions are possible for specific models. Coming up with a better numerical method that would allow one to get the relevant eigenvalues for truncation support cutoff larger than about $r=12$ sites, feasible with methods that we used, would be also a valuable extension. Note though that whether one has a prethermalization or DTC can be already inferred at small $r$, and so the method can be used as a simple tool to look for those regimes. In the kicked Ising model we in particular find a prethermalized DTC behavior for a wide range of parameters around $\tau=\pi$ (roughly $|\pi-\tau|<0.9$), with the effect though slowly decaying with increasing system size.
@@ A.1 The spectrum of $M(k)$
The size of matrix $M(k)$ is $\frac{3}{4}4^{r}$. For smaller $r$ the size is small enough to numerically diagonalize the whole matrix. For larger $r$ one can use methods targeting only the largest eigenvalue (in modulus), or few largest, like the simple power method for $\lambda_{1}$, or Lanczos based methods (we have used the ARPACK package arpack ) that construct a tridiagonal matrix representing $M$ on a Krylov subspace. The method does not require the whole matrix, which would be memory-wise prohibitive, but rather needs only application of $M$ on a vector. This can be done in-situ by simulating the action of our circuit on $r+2$ sites, see Eq.(9), that is, consecutively applying individual gates (Fig. 3).
@@
A.2 Correlation functions
We calculate correlation functions in finite systems numerically exactly by evolving a pure random initial state ${|\psi\rangle}$ in upto $L=33$ spins by repeatedly applying one and 2-qubit gates (see Fig. 3). For large $L$ such a single random state expectation value is a good approximation for an infinite temperature average. Namely, finite-size fluctuations in a correlation function scale as $\sim 1/2^{L/2}$, and are e.g. $\sim 10^{-5}$ for $L=32$. We always look at values of $C_{A}(t)$ that are larger than that (e.g. Fig. 5B). Note that in some cases, like studies of transport, there can be other boundary finite-size effects that are larger, see Fig. 7C for an example. Therefore we always check convergence with $L$, making sure to have finite-size effect under control. Memory requirements of the method are exponential in $L$ (keeping track of two pure states, each of size $N=2^{L}$), and reach $\approx 400\,$Gb for $L=33$, while each gate application takes ${\cal O}(N)$ operations.
@@ A.2 Correlation functions
In calculating the energy current autocorrelation function (19) in Fig. 7C we have simply used the kicked Ising model with a small timestep $\tau=0.05$. Comparing results at $L=32$ between $\tau=0.05$ and $\tau=0.02$ we can see that e.g. upto $t=30$ the difference in $D$ is about $0.002$ due to a finite Trotterization step approximation of Hamiltonian evolution. On the other hand, comparing different $L$ at $\tau=0.05$ we can estimate that the $L=33$ results do correctly reproduce the correlation function in the thermodynamic limit upto $t\approx 25$ with precision in $D$ of about $0.002$ (this time scales diffusively $\propto L^{2}$ with system size). Because $C_{J}(t)$ at $t=25$ is already quite small, taking into account that the remaining integral from $t=25$ to $\infty$ would bring only about $\approx 0.01$, we can estimate the diffusion constant by the integral of $L=33$ data upto $t=30$, which gives $1.452$, and estimate error at $0.01$.
One can in fact use the Green-Kubo integral upto a finite time $t$ upto which we trust to have the correct $C_{J}(t)$ in the TDL as a lower bound on $D$ because $C_{J}(t)$ for the chosen parameters is always positive, thereby estimating $D>1.45$. On the other hand, convergence of the truncated operator propagator method (Fig. 7A) seems to be monotonously from above, so using the $r=12$ value of $D=1.462$ we can estimate $D<1.46$. Both methods together therefore suggest tighter estimate $D=1.45-1.46$.
A.3 Boundary driven systems and NESS
@@
where $L_{j}$ are Lindblad operators. For our simulations we want to induce the flow of energy in the NESS and the Lindblad operators act on two boundary spins at each end. They are chosen in such a way that the dissipator ${\cal L}$ given solely by Lindblad operators on those two spins would induce a canonical state of two boundary spins. More precisely, $L_{j}$ at say the left edge are such that ${\cal L}(\rho_{L})=0$, where $\rho_{L}$ is the reduced density operator obtained by tracing over $4$ out of $6$ spins, $\rho_{L}\sim{\rm tr}({\rm e}^{-H/T_{L}})$, where $H$ is the Hamiltonian on $6$ spins. See Ref. jstat09 for more details. We use Lindblad (bath-like) “temperature” parameters $T_{L}=295$ and $T_{R}=300$, inducing an almost infinite-temperature NESS in which the average energy density is small.
@@ Appendix B BCH series
where $E_{3}\equiv[A,B]$, and we explicitly wrote out terms upto to 3rd order in products of $A$ and $B$. More systematically, the series can be written in such a way that all next order terms are a single commutator of two terms occurring in previous orders, i.e.,
@@
where $E_{j}\equiv[E_{j^{\prime}},E_{j^{\prime\prime}}]$ with appropriate $j^{\prime},j^{\prime\prime}<j$, starting with $E_{1}\equiv A,E_{2}\equiv B$ and $|1|=|2|=1$, while $z_{j}$ are rational numbers. Coefficients upto order $20$ are provided in Ref. casas08 ($|j|\leq 20$, i.e., $m\leq 19$ in Eq.26). For instance, the highest order $|j|=20$ contains $52377$ terms $E_{j}$, while all $20$ orders together have ${\rm max}(j)=111013$ terms.

### v2

@@ IV Conclusion
We have shown that the truncated propagator of extensive local operators is a very versatile method to study Floquet systems. It determines evolution of correlation functions, specifically, through the largest eigenvalue and the associated eigenvector it gives an explicit prediction for the prethermalization time and the effective (local) Floquet Hamiltonian governing prethermalization. This allowed us to resolve the question about reliability of approximating Hamiltonian evolution by a small-step Trotterization. Except for a single almost conserved observable all other orthogonal ones do not exhibit prethermalization, and even the almost conserved one decays in the thermodynamic limit. This means that due to the many-body dynamics the so-called shadowing time is always finite, and that there is no Trotterization transition in chaotic quantum systems. Previous observations of Trotterization transition were an artifact of finite-size effects. One has to ensure that the Heisenberg time is larger than the (long) prethermalization time. Because the limits $L\to\infty$ and $t\to\infty$ do not commute, in order to get correct thermodynamic results it is essential that our method works directly in the thermodynamic limit. If one would instead use small systems, for instance, exact diagonalization, one would get incorrect results.
@@
The truncated propagator can be also used to identify prethermal discrete time crystals. Just looking at the spectrum will already give an indication if there is a prethermal DTC (eigenvalues very close to roots of $1$, e.g., those around 1-1 in Fig. 2C at $\tau=3.0$). Studying chaotic kicked Ising model we find, perhaps surprisingly, that the prethermalized regime takes up the larger part of the gate duration interval $\tau\in[0,\pi]$. For instance, prethermalization is absent (say gap $\Delta>10^{-3}$, i.e. $T<10^{3}$) only for roughly $\tau\in[1.5,2.2]$. Glimpses of this rather slow thermalization in the 1D kicked Ising model can be seen already in data presented in Ref. prosen07 , and are present also in the 2D kicked Ising model. Prethermalization seems to be more common than previously thought and is not limited only to large frequencies (small $\tau$).
@@
Momentum dependence of the largest eigenvalue also contains information about transport. We have used it to calculate the energy diffusion constant during the prethermalization phase, as well as for the Hamiltonian model obtained in the limit $\tau\to 0$. Despite simply using small $\tau$ to approximate Hamiltonian dynamics, which causes the gap to be very small and difficult to calculate, the obtained diffusion constant is very precise.
@@
While we always used numerics to obtain the spectrum of the truncated propagator, it would be interesting to see if exact solutions are possible for specific models. Coming up with a better numerical method that would allow one to get the relevant eigenvalues for truncation support cutoff larger than about $r=14$ sites, feasible with the in-situ method that we use, would be also a valuable extension. Note though that whether one has a prethermalization or DTC can be already inferred at small $r$, and so the method can be used as a simple tool to look for those regimes. In the kicked Ising model we in particular find a prethermalized DTC behavior for a wide range of parameters around $\tau=\pi$ (roughly $|\pi-\tau|<0.9$), with the effect though slowly decaying with increasing system size.
We have also studied the kicked XX model, where prethermalization is found only for small $\tau$, as well as 2D kicked Ising model showing an interesting transition in the decay of correlation functions.
@@ A.1 The spectrum of $M(k)$
The size of matrix $M(k)$ is $N=\frac{3}{4}4^{r}$. For smaller $r$ the size is small enough to numerically diagonalize the whole matrix. For larger $r$ one can use methods targeting only the largest eigenvalue (in modulus), or few largest, like the simple power method for $\lambda_{1}$, or Lanczos based methods (we have used the ARPACK package arpack ) that construct a tridiagonal matrix representing $M$ on a Krylov subspace. The method does not require the whole matrix, which would be memory-wise prohibitive, but rather needs only application of $M$ on a vector. This can be simply done by simulating the action of our circuit on $r+2$ sites, see Eq.(9), that is, consecutively applying individual gates (Fig. 3). While the implementation on $r+2$ sites is straightforward, one can improve on speed and memory by directly performing evolution in-situ on just $r$ sites as described in the next subsection.
@@
A.2 In-situ propagation of operators
To get the action of the truncated propagator we need the overlaps of the propagated local operator $U^{\dagger}a^{(\bm{{\beta}})}_{1}U$, where $a^{(\bm{{\beta}})}_{1}$ lives on $r$ sites, with local operators that are either on the same $r$ sites, or are shifted by one site to the left or to the right (for evolutions that can shift by at most $1$ site, like the kicked Ising model), see Eq. (9). A simple way to do that is simply to evolve the operator $a^{(\bm{{\beta}})}_{1}$ on $r+2$ sites, starting by padding one identity at the left and the right edge, calculating the full $U^{\dagger}a^{(\bm{{\beta}})}_{1}U$ that now lives on $r+2$ sites, and then projecting it to the relevant operators with support on $r$ sites. However, that is expensive: we have to do evolution on $r+2$ sites and each additional site increases memory and computational requirements by a factor of $\approx 4$. Much faster is to do the in-situ evolution on just the relevant $r$ sites. In the following we describe how to do that, having in mind the 1d kicked Ising model.
One uses the fact that an initial operator with support on $r$ sites can increase its support only at its edges. Specifically, to get the contribution without any shift, $a^{(\bm{{\beta}})}_{1}\to a^{(\bm{{\alpha}})}_{1}$, the two qubit gate acting on the boundary (dashed lines in Fig. 12(a)) has to map at the left edge $\mathbbm{1}_{0}x_{1}$, where “xx” indicates any non-identity operator, $x\in\{\sigma^{\rm x},\sigma^{\rm y},\sigma^{\rm z}\}$, to again the identity at the leftmost site (see Fig. 12(a)). For the kicked Ising model where the 2-qubit gate has only the $zz$ interaction the only possibility is $\mathbbm{1}_{0}x_{1}\to\mathbbm{1}_{0}x_{1}$. At the right edge one has a similar situation, with the only difference that the last operator can be also the identity, resulting in $(*)_{r}\mathbbm{1}_{r+1}\to(*)_{r}\mathbbm{1}_{r+1}$, where $(*)\in\{\sigma^{\rm x},\sigma^{\rm y},\sigma^{\rm z},\mathbbm{1}\}$.
Figure 12: Truncated propagator construction: in-situ propagation of local operators on just $r$ sites. The initial operator is supported on $r$ sites (black circles) padded with an identity outside its support (white circles). Executing a 2-qubit $zz$ gate on dashed lines has to execute a mapping indicated by the arrow. Afterwards one has to perform the remaining 2-qubit gates (full lines). See text for details.
If we want to achieve a mapping of the initial operator to a shifted one supported on $r$ consecutive sites we have to get an identity at its tail end, and a non-identity operator at its front-end. For the $zz$ gate in question such a mapping can be achieved only by $\mathbbm{1}_{0}y_{1}\to\sigma^{\rm z}_{0}\bar{y}_{1}$, where “yy” stands for $y\in\{\sigma^{\rm x},\sigma^{\rm y}\}$, and its negation $\bar{y}$ is $\sigma^{\rm x}$ if $y=\sigma^{\rm y}$ and $\sigma^{\rm y}$ if $y=\sigma^{\rm x}$. This is sketched in Fig. 12(b). For a shift to the left there is another possibility, namely, the right-most site can already have an identity operator, resulting in a scheme in Fig. 12(c) (for a shift to the right (b) is the only option). We can see that in both Fig. 12(b-c) it is enough to keep evidence of only $r$ sites – at the beginning on $r$ black circles, and after performing the first step, transformation on all dashed links, only on the leftmost $r$ sites, afterwards executing the remaining gates (full-line links in Figure).
A.3 Correlation functions
We calculate correlation functions in finite systems numerically exactly by evolving a pure random initial state $|\psi\rangle$} in up-to $L=33$ spins by repeatedly applying one and 2-qubit gates (see Fig. 3). For large $L$ such a single random state expectation value is a good approximation for an infinite temperature average. Namely, finite-size fluctuations in a correlation function scale as $\sim 1/2^{L/2}$, and are e.g. $\sim 10^{-5}$ for $L=32$. We always look at values of $C_{A}(t)$ that are larger than that (e.g. Fig. 5B). Note that in some cases, like studies of transport, there can be other boundary finite-size effects that are larger, see Fig. 7C for an example. Therefore we always check convergence with $L$, making sure to have finite-size effect under control. Memory requirements of the method are exponential in $L$ (keeping track of two pure states, each of size ${\cal N}=2^{L}$), and reach $\approx 400\,$Gb for $L=33$, while each gate application takes ${\cal O}({\cal N})$ operations.
@@ A.3 Correlation functions
In calculating the energy current autocorrelation function (19) in Fig. 7C we have simply used the kicked Ising model with a small timestep $\tau=0.05$. Comparing results at $L=32$ between $\tau=0.05$ and $\tau=0.02$ we can see that e.g. up-to $t=30$ the difference in $D$ is about $0.002$ due to a finite Trotterization step approximation of Hamiltonian evolution. On the other hand, comparing different $L$ at $\tau=0.05$ we can estimate that the $L=33$ results do correctly reproduce the correlation function in the thermodynamic limit up-to $t\approx 25$ with precision in $D$ of about $0.002$ (this time scales diffusively $\propto L^{2}$ with system size). Because $C_{J}(t)$ at $t=25$ is already quite small, taking into account that the remaining integral from $t=25$ to $\infty$ would bring only about $\approx 0.01$, we can estimate the diffusion constant by the integral of $L=33$ data up-to $t=30$, which gives $1.452$, and estimate error at $0.01$.
One can in fact use the Green-Kubo integral up-to a finite time $t$ up-to which we trust to have the correct $C_{J}(t)$ in the TDL as a lower bound on $D$ because $C_{J}(t)$ for the chosen parameters is always positive, thereby estimating $D>1.45$. On the other hand, convergence of the truncated operator propagator method (Fig. 7A) seems to be monotonously from above, so using the $r=12$ value of $D=1.462$ we can estimate $D<1.46$. Both methods together therefore suggest tighter estimate D=1.451.46D=1.45-1.46.
A.4 Boundary driven systems and NESS
@@
[…]
@@ Appendix B BCH series
where $E_{3}\equiv[A,B]$, and we explicitly wrote out terms up-to to 3rd order in products of $A$ and $B$. More systematically, the series can be written in such a way that all next order terms are a single commutator of two terms occurring in previous orders, i.e.,
@@
where $E_{j}\equiv[E_{j^{\prime}},E_{j^{\prime\prime}}]$ with appropriate $j^{\prime},j^{\prime\prime}<j$, starting with $E_{1}\equiv A,E_{2}\equiv B$ and $|1|=|2|=1$, while $z_{j}$ are rational numbers. Coefficients up-to order $20$ are provided in Ref. casas08 ($|j|\leq 20$, i.e., $m\leq 19$ in Eq.27). For instance, the highest order $|j|=20$ contains $52377$ terms $E_{j}$, while all $20$ orders together have ${\rm max}(j)=111013$ terms.
@@
Figure 16: 1D kicked XX model (31). (A) shows the circuit, (B) the spectra of truncated propagators for $r=7$, (C) autocorrelation function of magnetization (black dashed curves use the gap from (D)), and (D) the gap for $r=8$. Two crosses indicate the decay rate for dashed curves for $\tau$ in (C).